OpenStudy (anonymous):

f(x)=xsquared+6 and g(x)=x+8/8 Find ( g o f)(-7)

OpenStudy (bahrom7893):

this just means g(f(-7))

OpenStudy (bahrom7893):

g(f(x) = (x^2+6 + 8)/8 g(f(-7)) = (49 + 14)/8